Oslo Company prepared the following contribution format income statement based on a sales volume of 1,000 units (the relevant range of production is 500 units to 1,500 units):
Sales $ 20,600
Variable expenses 12,200
Contribution margin 8,400
Fixed expenses 6,468
Net operating income $ 1,932
1.If the variable cost per unit increases by $1.20, spending on advertising increases by $1,700, and unit sales increase by 250 units, what would be the net operating income? (Do not round intermediate calculations.)

2. What is the break-even point in unit sales? (Do not round intermediate calculations.)

3. What is the break-even point in dollar sales? (Round intermediate calculations to 4 decimal places. Round your answer to the nearest dollar amount.)

4.
A. What is the margin of safety percentage? (Round your final answers to the nearest whole percentage (i.e, .12 should be entered as 12)
B. What is the margin of safety in dollars? (Do not round intermediate calculations.)
Answer #1

Currently

Volume = 1000

Sales = $20600

Unit price = $20.6

Variable expense = $12200

Unit variable cost = $12.2

Fixed expense = $6468

Net operating income = 20600 – 12200 – 6468 = 1932

1

New variable cost = 12.2 + 1.2 = $13.4

New volume = 1250

New fixed cost = 6468+1700 = $8168

Net operating income = 1250*20.6 – 1250*13.4 – 8168 = $832

(Assuming that the following questions 2, 3, 4, 5 are based on the data from question 1 and not the current state)

2

Break-even point = Fixed cost / (Unit price – Variable cost)

BEP = 8168/(20.6-13.4) = 1134.4 or 1135 units

3

Break-even point in dollar sales = Fixed cost / Contribution margin ratio

Contribution margin ratio = Contribution margin/ Sales = (20.6-13.4)/20.6 = 0.349

BEP in dollar sales = 8168/0.349 = $23369.5

4

Margin of safety = (Sales level – BEP)/Sales level

MoS = (1250-1135)/1250 = 0.092 or 9.2%

5

MoS in dollars = 1250*20.6 – 23369.5 = 2380.5

We have assumed that the number 2, 3, 4 and 5 are based on the information modified in question 1. In case that is not, then use the original values and you will get the answer.